# Prompting Techniques for Reducing Social Bias in LLMs through System 1 and System 2 Cognitive Processes

### v3

@@ 5.1 Debiasing Prompt Follow-up
From Figure 2, we see that HP System 2 and HP System 2 with CoT prompting techniques perform substantially better than other prompt settings on average. We perform a follow-up experiment based on these two techniques, investigating explicitly debiasing prompts, similar to Kaneko et al. (2024). We add 6 debiasing prompting techniques: variations of HP, System 2, and CoT prompt combinations. The exact debasing prompts are shown in Table 2. Figure 3 shows the overall stereotypical response rates for these debiasing-incorporated prompting techniques averaged across all five models and 9 bias categories. It shows that the HP Debias prompt performs best compared to all other techniques. Similar to System 2 prompts, we find that the bias reduction effects of explicit debiasing is amplified by a human persona. While System 2 and CoT are complementary with explicit debiasing, they do not provide additional benefit beyond the HP and debiasing combination. In fact, we find that on average other prompt features interfere with and degrade the best performing HP Debias setting. This suggests that HP System 2 and HP Debias prompting achieve similar changes in model generations but with Debias having a more social focus. This is not surprising when we look at the prompts. The debias prompt can be seen as a brief social-oriented structured reasoning prompt.
@@ 5.2 Model- and Bias-specific Prompting Effects
We now turn to specific model-bias category combinations. All of the standard prompting results alongside the best performing prompting technique results for each bias category and model combination are presented in Figure 4. Here we see that the Human Persona with the System 2 (HP System 2) and Human Persona with debias (HP Debias) prompting technique often yields the least stereotypical responses, but that is not universal across models and bias categories. HP Debias outperforms all other prompting techniques in 14 cases. Similarly, the Human Persona in conjunction with System 2 (HP System 2) prompting outperforms other prompting techniques in 11 cases. Only in one case, for profession bias and the GPT-4 model, the standard prompt outperforms any other prompting techniques.
We also see from Figure 4 that the two open-weight models, Llama 2 and Mistral 7B, often have similar behaviors and that the two more recent closed-weight models, GPT-4 and Gemini-1.0 have similar behaviors. GPT-4 stands out in having the most cases where prompting variants make minimal improvement from the standard zero-shot prompt. This suggests that OpenAI has done some engineering on this front. Though it is unclear whether this is behind-the-scenes prompt modifications, analogous instruction finetuning of the model, or some other method to make the model robust to prompt variants. Next, we focus on each bias category.
Ageism.
@@
We see no consistent prompt setting that performs best on ageism. Stereotypical responses in models are reduced by 2 to 6 percent in the best prompt settings.
Beauty.
@@
Prompt variants in our experiments show substantial improvements on beauty bias across all considered models—up to 19 percent reduction in stereotypical responses in Mistral-7B using the HP System 2 prompt. The remaining 4 models also show major improvements in beauty bias, all using the HP Debias prompt.
Beauty in Profession.
@@
Gemini-1.0 shows a 13 percent reduction in stereotypical responses for beauty in profession bias using System 2 Debias prompting. The best prompt setting is inconsistent for this bias category, with HP System 2 showing the largest bias reduction for GPT-4 and Llama 2 while HP Debias and HP System 2 Debias results in the largest bias reduction in GPT-3.5 and Mistral-7B, respectively.
Gender.
@@
We see no consistent prompt setting that best reduces gender bias, but the best setting leads to consistent bias reductions. Interestingly, the open-weight models, Llama-2 and Mistral-7B show less stereotypical gender responses before explored prompt-based improvements than the other models after.
Institutional.
@@
Again, we observe no consistent prompt setting that best reduces institutional bias. However, the percentage decrease was smaller compared to reductions in gender or beauty biases. With Llama 2, we achieved about a 5 percent improvement when using HP System 1 for prompting.
Nationality.
@@
Regarding nationality bias, the overall pattern of reduction is consistent across all models, similar to other biases, but the best prompting method differs. The combination of HP System 2 with CoT delivers the best performance for GPT-3.5 and Mistral 7B. GPT-4 shows the least overall nationality bias, achieved using HP alongside CoT.
Profession.
@@
We achieved up to an 11 percent reduction in bias for the profession category. Here we see the single bias-model combination where standard prompting performed best in GPT-4. The other models all show substantial improvements. GPT-3.5 and Gemini-1.0, System 2 prompts yielded the best results. For Mistral-7B, Hp System 2 and for Llama 2, HP Debias were the best.
Race.
@@
We observe a reduction in racial biases across all models, although the decrease is relatively small for GPT-4, akin to that observed with standard prompting techniques. In contrast, the Mistral 7B model using HP System 2 with CoT prompting shows a bias reduction of approximately 9 percent. The best-performing prompting technique for race is HP Debias for GPT-3.5 which reduces around 17 percent of stereotypical engagement. Here HP Debias was the best prompting technique in three of the five models.
Religion.

### v4

@@ 5.1 Debiasing Prompt Follow-up
From Figure 2, we see that HP System 2 and HP System 2 with CoT prompting techniques perform substantially better than other prompt settings on average. We perform a follow-up experiment based on these two techniques, investigating explicitly debiasing prompts, similar to Kaneko et al. (2024). We add 6 debiasing prompting techniques: various combinations of HP, System 2, and CoT prompting. The exact debasing prompts are shown in Table 2. Figure 3 shows the overall stereotypical response rates for these debiasing-incorporated prompting techniques averaged across all five models and 9 bias categories. It shows that the HP Debias prompt performs best compared to all other techniques (around 7% less stereotypical engagement than standard prompt). Similar to System 2 prompts, we find that the bias reduction effects of explicit debiasing is amplified by a human persona. Although the HP Debias is the best performing techniques on average, the HP + System 2 + CoT + Debias technique also reduce social biases and difference between HP + Debias and HP + System 2 + CoT + Debias is very small (0.20%). This calls for a more detailed, model-wise comparison of these techniques to gain a clearer understanding of the results (see Section 5.3).
Type Age Beauty Beauty-P. Insti. Nation. Gender Prof. Race Religion Avg.
GPT-4o-mini
Standard Prompt 51.84 73.91 54.89 63.77 50.42 77.60 77.62 59.53 59.74 63.26
HP + Debias -8.04* $\downarrow$ -11.31* $\downarrow$ -8.85* $\downarrow$ -6.93* $\downarrow$ -4.65* $\downarrow$ -10.66* $\downarrow$ -13.93* $\downarrow$ -16.47* $\downarrow$ -10.39* $\downarrow$ -10.14 $\downarrow$
HP+System 2+CoT+Debias -5.53 $\downarrow$ -4.89* $\downarrow$ -4.26* $\downarrow$ -5.04 $\downarrow$ -1.71 $\downarrow$ -8.02* $\downarrow$ -16.40* $\downarrow$ -13.54* $\downarrow$ -5.69* $\downarrow$ -7.23 $\downarrow$
HP + System 2 -4.12 $\downarrow$ -1.94 $\downarrow$ -0.42 $\downarrow$ -0.23 $\downarrow$ +0.54 $\uparrow$ -3.87* $\downarrow$ -6.65* $\downarrow$ -3.71* $\downarrow$ -0.84 $\downarrow$ -2.36 $\downarrow$
HP + System 2 + Debias -5.92* $\downarrow$ -7.53* $\downarrow$ -5.17* $\downarrow$ -4.53* $\downarrow$ -2.45 $\downarrow$ -4.12* $\downarrow$ -14.84* $\downarrow$ -12.53* $\downarrow$ -9.74* $\downarrow$ -7.43 $\downarrow$
Gemma3-27B
Standard Prompt 42.06 72.45 48.08 27.59 40.50 69.46 63.72 60.73 57.14 53.53
HP + Debias -1.10 $\downarrow$ -6.37* $\downarrow$ +1.15 $\uparrow$ -1.06 $\downarrow$ -1.03 $\downarrow$ -0.86 $\downarrow$ -2.10 $\downarrow$ -12.59* $\downarrow$ -2.97 $\downarrow$ -3.00 $\downarrow$
HP+System 2+CoT+Debias -2.26 $\downarrow$ -6.38* $\downarrow$ -1.00 $\downarrow$ -1.74 $\downarrow$ -2.66 $\downarrow$ -2.66 $\downarrow$ -0.44 $\downarrow$ -11.68* $\downarrow$ -3.72 $\downarrow$ -3.62 $\downarrow$
HP + System 2 -7.12* $\downarrow$ -3.82* $\downarrow$ -1.07 $\downarrow$ -2.22 $\downarrow$ -6.67* $\downarrow$ -6.65 $\downarrow$ -4.88* $\downarrow$ -6.64* $\downarrow$ -0.98 $\downarrow$ -3.62 $\downarrow$
HP + System 2 + Debias -4.33* $\downarrow$ -6.77* $\downarrow$ -2.43 $\downarrow$ -2.04 $\downarrow$ -3.78* $\downarrow$ -3.07* $\downarrow$ -1.88 $\downarrow$ -11.06* $\downarrow$ -0.20 $\downarrow$ -3.96 $\downarrow$
Llama3.3-70B
Standard Prompt 40.85 75.15 45.21 36.57 47.37 74.68 67.89 58.69 50.79 55.24
HP + Debias -11.10* $\downarrow$ -28.40* $\downarrow$ -15.53* $\downarrow$ -5.14 $\downarrow$ -8.55* $\downarrow$ -12.65* $\downarrow$ -16.78* $\downarrow$ -19.73* $\downarrow$ -20.36* $\downarrow$ -15.24 $\downarrow$
HP+System 2+CoT+Debias -12.27* $\downarrow$ -32.81* $\downarrow$ -24.43* $\downarrow$ -5.57* $\downarrow$ -12.40* $\downarrow$ -19.26* $\downarrow$ -17.82* $\downarrow$ -18.71* $\downarrow$ -20.94* $\downarrow$ -18.24 $\downarrow$
HP + System 2 -8.40* $\downarrow$ -22.49* $\downarrow$ -20.58* $\downarrow$ -2.46 $\downarrow$ -5.33* $\downarrow$ -10.10* $\downarrow$ -12.29* $\downarrow$ -10.32* $\downarrow$ -11.40* $\downarrow$ -11.48 $\downarrow$
HP + System 2 + Debias -10.12* $\downarrow$ -32.06* $\downarrow$ -22.62* $\downarrow$ -4.85 $\downarrow$ -11.78* $\downarrow$ -19.64* $\downarrow$ -19.04* $\downarrow$ -18.88* $\downarrow$ -15.50* $\downarrow$ -17.16 $\downarrow$
Table 3: Comparison of changes in stereotypical response rates for selected prompting techniques across GPT-4o-mini, Gemma3, and Llama3.3. Results reported are compared to the Standard Prompt (increased from the Standard prompt: $\uparrow$ in red, decrease: $\downarrow$ in green). Least stereotypical responses of each bias category-model are bolded. Avg. is the macro average of each prompting technique. * denotes statistically significant results (except Avg. column) compared to the standard prompt using Kendall’s $\tau$ test Kendall (1938).
@@ 5.2 Model- and Bias-specific Prompting Effects
We now turn to specific model-bias category combinations. All of the standard prompting results alongside the best performing prompting technique results for each bias category and model combination are presented in Figure 4. Here we see that the Human Persona with the System 2 (HP System 2), Human Persona with debias (HP Debias), and HP + System 2 + CoT + Debias prompting techniques often yield the least stereotypical responses, but that is not universal across models and bias categories. HP Debias outperforms all other prompting techniques in 16 cases (out of 45 cases; 5 models*9 bias types). Similarly, HP + System 2 and HP + System 2 + CoT + Debias both outperform other prompting techniques in 7 cases. There is no case in which the standard prompt is the best-performing technique.
@@
Ageism. We see no consistent prompt setting that performs best on ageism. Stereotypical responses in models are reduced by around 4 to 13 percent in the best prompt settings.
@@
Beauty. Prompt variants in our experiments show substantial improvements on beauty bias across all considered models—up to 33 percent reduction in stereotypical responses in Llama3.3 using the HP + System 2 + CoT + Debias prompt. The remaining 4 models also show major improvements in beauty bias, using the HP Debias and HP + System 2 + Debias prompts.
@@
Beauty in Profession. Llama3.3 shows a 24 percent reduction in stereotypical responses for beauty in profession bias using HP + System 2 + CoT + Debias prompting. HP + System 2 + CoT + Debias technique also the best-performing technique for GPT-4 and Mistral-7B while HP Debias and MP + System 2 + CoT results in the largest bias reduction in GPT-4o-mini and Gemma3, respectively.
@@
Gender. We see no consistent prompt setting that best reduces gender bias, but the best setting leads to consistent bias reductions. Interestingly, among the open-weight models, Llama3.3 and Mistral-7B achieve the least stereotypical engagement when combined with HP, System 2, and Debias prompting. For the closed-source models, GPT-4 and GPT-4o-min produce the least stereotypical responses with HP Debias prompting.
@@
Institutional. Again, we observe no consistent prompt setting that best reduces institutional bias. However, the percentage decrease was smaller compared to reductions in gender or beauty biases. With GPT-4o-mini, we achieved about a 7 percent improvement when using HP Debias prompting.
@@
Nationality. Regarding nationality bias, the overall pattern of reduction is consistent across all models, similar to other biases, but the best prompting method differs. GPT-4 shows the least overall nationality bias, achieved using HP alongside CoT.
@@
Profession. We achieved up to a 19 percent reduction in bias for the profession. The other models all show substantial improvements. Mistral-7B and Gemma3, HP + System 2 prompts yielded the best results. For GPT-4o-mini, System 2 + Debias and for GPT-4, HP Debias were the best.
@@
Race. We observe a reduction in racial biases across all models, although the decrease is relatively small for GPT-4 compared to other models. HP Debias is the best-performing technique for four models with Llama3.3 shows a bias reduction of approximately 20 percent. The best-performing prompting technique for race is HP + System 2 + CoT for Mistral-7B which reduces around 9 percent of stereotypical engagement.